OpenStudy (mackenzie2013):

What is the solution to the equation x/3 + 5= 8? x= 1 x= 6 x= 7 x= 9

OpenStudy (anonymous):

x/3 + 5= 8 x/3=8-5 x/3=3 x=3x3 x=9 So, it's option number D :)

OpenStudy (anonymous):

I hope I'm clear :)

OpenStudy (mackenzie2013):

Thanks !